# Project Page: https://gyhdog99.github.io/projects/mocle/

### v4

@@ 3.5 Universal Expert
However, as shown in Sec. 4.4.1, we find that the formulation in Sec. 3.4 still hurts the generalization ability of the entire model. We hypothesize that this might be attributed to the fact that the instruction-tuned models generalize to unseen tasks by training on extensive instructions [60], while in our formulation, each expert sees fewer instructions than the original dense model. Hence, we propose a universal expert that could learn from all the training data. Specifically, we fix the number of activated experts to 1 (i.e., $k$ equals 1) and define $G_{\text{max}}$ as the maximum element in $\mathbf{G}$. Then the output for all the experts is expressed as,
@@
In this formulation, we additionally train a universal expert parameterized by $\mathbf{W}_{u}$. Different from task experts that are activated only for specific model inputs, the universal expert is activated for any inputs. The final output is the weighted sum of outputs from one of the experts and the universal expert puls the original LVLM’s output. The weighting for this universal expert is the complement of that of the chosen task expert. Consequently, the task expert learns distinct skills for certain tasks while the universal expert masters holistic understanding of the training corpus. The synergy between them offers both specialization and generalization for the LVLMs with MoCLE.
@@ 4 Experiment
In this section, we conduct an assessment of MoCLE across multiple downstream tasks under a zero-shot setting. Initially, we detail the experimental settings and implementation details, which is followed by a description of the datasets and instructions employed, along with the outcomes of our evaluations. Lastly, we present an ablation study and visualizations of clustering and routing results.
4.1 Settings and Implementation Details
In all the experiments, we utilize the InstructBLIP[14] model, particularly its Vicuna-7B variant [70], as the base LVLM to assess the effectiveness of MoCLE. We encode all the instructions of different datasets using the all-MiniLM-L6-v2 variant of the Sentence Transformer model [49] and cluster their embeddings into $64$ clusters via $k$-means clustering algorithm. We apply the proposed MoCLE on the LLM (i.e., Vicuna-7B) of InstructBLIP. Specifically, for each transformer layer in the LLM, we insert the MoCLE, which consists of 4 task LoRA experts and one LoRA universal expert, into the query and value projection of self-attention blocks. Inside each expert, we set the rank of the LoRA to 8 and the temperature $\tau$ to $0.05$. During instruction tuning, we only finetune parameters of Q-Former and the LoRA experts while keeping other parts of the LLM frozen.
Models GQA VSR IconQA Visdial MME POPE A-OKVQA OKVQA VQAv2 Direct MC (test) (test-dev) LLaVA (13B) 41.3 51.2 43.0 - - 67.0 50.5 - 54.4 - MiniGPT-4 (13B) 30.8 41.6 37.6 - 581.7 71.4 38.7 - 37.5 - InstructBLIP (7B) 48.6 60.8 43.4 46.3 1202.9 77.6 58.8 73.8 57.0 77.4 49.3 64.7 46.3 46.9 1222.6 82.1 61.5 78.2 59.8 78.9 MoCLE (7B) (+0.7) (+3.9) (+2.9) (+0.6) (+19.7) (+4.5) (+2.7) (+4.4) (+2.8) (+1.5)
Table 1: Zero-shot results on the held-out datasets (GQA, VSR, IconQA, Visdial) and LVLM benchmarks (MME, POPE) and evaluation on held-in datasets (A-OKVQA, OKVQA, VQAv2). Here Direct and MC denote directly answering and multiple choices. Best results are marked in bold.
Models Flickr 30K TextVQA HM SciQA image MSVD QA MSRVTT QA iVQA InstructBLIP (7B) 81.3 53.9 65.3 62.0 41.4 23.0 51.3 81.9 57.1 65.6 63.9 42.6 24.4 53.2 MoCLE (7B) (+0.6) (+3.2) (+0.3) (+1.9) (+1.2) (+1.4) (+1.9)
Table 2: Zero-shot results on the held-out datasets. Here, HM and SciQA denote HatefulMemes and ScienceQA, respectively.
4.2 Datasets and Instructions
To demonstrate the effectiveness of MoCLE, we follow InstructBLIP to use the same training and evaluation data unless mentioned explicitly. As Web CapFilt[28], one of the training datasets, is not fully available, we can only download a subset of them. Besides this, we utilize all the other training data, including COCO [35], TextCaps [55], VQAv2 [19], OKVQA [45], A-OKVQA [52], OCR-VQA [46], and LLaVA-Instruct150K [38], for instruction tuning.
For the zero-shot evaluation datasets, we follow InstructBLIP but only keep Flickr30K [64], GQA [22], VSR [36], IconQA [43], TextVQA [56], Hateful Memes [25], ScienceQA [42], MSVD-QA [61], MSRVTT-QA [61] and iVQA [62]. NoCaps dataset [1] is originally utilized for the image captioning task, which, however, is not adopted here since we cannot reproduce the reported results, probably due to the incomplete Web CapFilt dataset used. In addition, we consider another two widely adopted LVLM benchmarks, i.e., MME [16] and POPE [33] for zero-shot evaluations. For held-in evaluation, we consider the VQAv2 [19], OKVQA [45] and A-OKVQA [52].
For ScienceQA, we only evaluate the set with image context. Following previous works [14], we report the CIDEr score [57] for Flickr30K, the iVQA accuracy for iVQA, AUC score for HatefulMemes, and Mean Reciprocal Rank (MRR) for Visual Dialog. For LVLM benchmarks, we report the perception score for MME and F1 score for the adversarial split of POPE. For all other datasets, we report the top-1 accuracy (%). The instructions to conduct all the experiments in this paper are from the original InstructBLIP paper. The detailed introductions of the held-in and held-out tasks/datasets are also presented in [14].
@@ 4.3 Evaluation Results
Table 1 and Table 2 present the experimental results on multiple zero-shot and held-in vision-language tasks. We build our method upon the Vicuna-7B variant of InstructBLIP. Compared to this model, the proposed MoCLE method shows considerable performance improvement. Specifically, on held-out datasets such as IconQA, Visual Spatial Reasoning (VSR), TextVQA and ScienceQA datasets, we obtain an absolute performance gain of 2.9% 3.9%, 3.2%, and 1.9%, respectively. This indicates that the proposed MoCLE method facilitates generalization to unseen tasks/instructions since these datasets are unseen during training.
In addition, we observe considerable improvement on common LVLM benchmarks (i.e., MME and POPE). POPE evaluates whether the model suffers from object hallucinations while MME considers a wide range of abilities such as math, OCR, color and understanding of arts. Therefore, our MoCLE cannot only promote task generation, but also improve overall abilities of the LVLM.
Further, the held-in tasks whose datasets are seen during training could also benefit from the proposed MoCLE. Specifically, with our MoCLE, an absolute improvement of 4.4%, 2.8% and 1.5% could be observed on A-OKVQA (MC), OKVQA and VQAv2. This indicates that MoCLE could effectively alleviate the task conflicts during multi-task learning.
@@ 4.4 Ablation Studies
In this section, we first ablate the effectiveness of the main components (e.g., task experts and universal expert) in the proposed MoCLE method. Then we conduct a thorough analysis to study how the proposed MoCLE method would respond to the change of hyper-parameters (e.g., temperature and the number of clusters and task experts). Notice that for the ablations, we use Flickr30K (Flickr), Hateful Memes (HM), ScienceQA (SciQA), IconQA, Visual Spatial Reasoning (VSR) and TextVQA for evaluation.
Components Flickr HM SciQA IconQA VSR TextVQA Avg.
InstructBLIP 81.3 65.3 62.0 43.4 60.8 53.9 61.1
+LoRA (rank=8) 81.3 65.1 57.4 44.2 62.8 49.4 60.0
 (rank=64) 81.5 65.2 62.0 43.9 62.6 49.0 60.7
+Mixture (Cluster) 81.9 65.4 63.3 46.1 58.9 54.9 61.8
+Universal Expert 81.9 65.6 63.9 46.3 64.7 57.1 63.3
Table 3: Comparison of individual components of the MoCLE framework in zero-shot vision-language tasks. We include three ablation studies that cumulatively add (i) parameter-efficient finetuning of LLM using LoRA of different ranks (higher ranks are ablated, i.e. rank=64, but we use rank=8 for ablations and MoCLE), (ii) replacing the single LoRA with its mixture and route inputs by their cluster, and (iii) training with the universal expert. Default settings are marked in gray.
4.4.1 Effects of Different Components.
We start from vanilla InstructBLIP with Vicuna-7B and add the main components in the proposed MoCLE method step by step. As shown in Table 3, naïvely applying LoRA to InstructBLIP could be problematic. Although we observe minor performance gain in IconQA and VSR, the peformance on ScienceQA and TextVQA are much inferior to InstructBLIP. We speculate that a single LoRA offers very limited flexibility, thus it performs not so well on many tasks. Note that we further conduct an experiment with a rank of 64 (much larger than $r=8$ in the default setting of MoCLE). However, the resulting model is still outperformed by InstructBLIP with less trainable parameters. This indicates that naïvely increasing the capacity of the network cannot mitigate the task conflicts in the training data. However, when we replace a single LoRA with the Mixture of LoRAs and route the input based on the cluster of their instructions, the resulting model outperforms its predecessor by a large margin on SciQA, IconQA, and TextVQA. Nevertheless, we notice a sharp performance drop in the VSR task. This could be due to the fact that instruction-tuned model generalizes to unseen tasks by training on many instructions, while in our case, each expert sees fewer instructions than the original dense model does. As shown in Table 3, this issue could be alleviated by the usage of the universal expert, which is updated by the gradients from all the training data. The same result can be observed in TextVQA, which requires not only the VQA ability but also additional optical character recognition (OCR) skills. Specifically, in the training stage, we have instructions “Question: {}. Short answer:” for all VQA datasets and “OCR tokens: {}. Write a description for the photo” for TextCaps [55], while the TextVQA [56] instruction is “OCR tokens: {}, Question: {}. Short answer:”, requiring LVLMs to combine knowledge from both tasks.
4.4.2 Universal Expert vs. Top-2 Experts.
MoCLE shares similarities with the top-2 variant of MoE. Specifically, they both activate two experts each time. However, the universal expert differs from top-2 MoE by consistently selecting the universal expert, whose contribution is the complement of the top-1 expert. As shown in Table 4, the top-2 MoE model yields inferior results compared to MoCLE. Therefore the naïvely activating one more expert cannot address the generalization problem. Surprisingly, this variant performs even worse than the MoCLE model without the universal expert reported in Table 3. This might be resulted from task conflicts with more parameters shared.
Flickr HM SciQA IconQA VSR TextVQA Avg.
@@
Gating strategy Flickr HM SciQA IconQA VSR TextVQA Avg.
Token 81.7 65.4 61.9 44.0 49.0 46.6 58.1
Sentence 82.0 65.1 62.3 45.3 56.6 47.0 59.7
Dataset 80.3 64.6 63.1 45.9 62.6 55.8 62.1
Cluster 81.9 65.6 63.9 46.3 64.7 57.1 63.3
@@
Table 5: Ablation study on routing inputs based on different conditions. Default settings are marked in gray.
4.4.3 Gating Strategies.
@@
We compare the proposed cluster-conditioned gating strategy with Token-MoLE, Sentence-MoLE and Dataset-MoLE, where MoLE denotes the mixture of LoRA experts by applying the MoE technique to LoRA, Token-MoLE obtains the routing decision based on the hidden representations of each token, and Sentence-MoLE averages the representations of the instruction tokens while excluding the visual tokens. Dataset-MoLE is a special case of MoCLE as it treats each dataset as a cluster while MoCLE leverages $k$-means to obtain clusters among all the datasets. Note that the universal expert is adopted for all these MoLEs. As observed in Table 5, Token-MoLE and Sentence-MoLE give inferior results on the evaluation tasks. We speculate that this is because (1) an expert only learns a subset of the tasks, giving poor generalization and (2) similar tasks are not grouped together by the same expert, resulting in task conflicts within that expert. The later can be evidenced by the routing visualization in Figure 4, where samples in the same dataset are routed to multiple experts instead of a dedicate one. Further, we observe performance drops when replacing $k$-means clustering with statically treating each dataset as a cluster. This results from the fact that Dataset-MoLE is less flexible than MoCLE because it can only assign a dataset to one cluster. However, in practice, we observe multiple tasks in a dataset which should be assigned to different clusters. (e.g., llava_150k contains reasoning, conversations and captioning, which are assigned to different clusters/experts as in Figure 2(a) and 3(a)).
Table 6: Ablation study on the number of clusters $K$.
# Clusters Flickr HM SciQA IconQA VSR TextVQA
4 81.3 65.2 63.4 45.3 58.3 55.3
16 82.0 64.3 61.7 46.3 67.2 56.9
64 81.9 65.6 63.9 46.3 64.7 57.1
256 82.5 65.6 63.2 46.0 57.0 56.9
Temperature $\tau$ Flickr HM SciQA IconQA VSR TextVQA
0.01 81.6 64.7 61.7 46.3 61.2 55.8
0.05 81.9 65.6 63.9 46.3 64.7 57.1
0.1 82.3 65.7 63.7 45.7 63.0 56.7
0.2 81.9 64.8 62.1 45.2 61.3 54.0
Table 6: Ablation study on the number of clusters $K$.
Table 7: Ablation study on the temperature $\tau$ in Eq. (1).
#Task experts Flickr HM SciQA IconQA VSR TextVQA
2 81.4 64.4 62.8 45.6 58.1 50.9
4 81.9 65.6 63.9 46.3 64.7 57.1
8 81.0 65.3 63.5 45.9 65.0 57.4
Table 8: Ablation study on the number of task experts.
4.4.4 Number of Clusters.
@@
The merit of sentence encoding and clustering lies in reducing the complexity of the instruction, offering robust and semantic-relevant routing process. The number of clusters $K$ in this process could measure such complexity. A extremely small $K$ would diminish the difference of various tasks, result in many different tasks to be processed by the same expert. This could increase the chance of task conflicts. As shown in Table 7, when we cluster the inputs into 4 groups, the resulting model performs poorly on the evaluation tasks. However, as we increase the number of clusters to 16 and 64, we observe considerable performance gains. Nevertheless, too large $K$ introduces unnecessary complexity in the routing process, e.g., a paraphrased instruction gets routed to different experts. So we use 64 clusters as the default setting.
4.4.5 Number of Task Experts.
As demonstrated in Table 8, more task experts usually provides with stronger capacity. Specifically, when only 2 task experts are employed, we observe inferior results on VSR and TextVQA. This model has similar capacity to the single LoRA model in Sec. 4.4.1, where only one LoRA encounters difficulties in fitting a diverse set of tasks. When the number of task experts is increased to 4, the performance gets improved. Specifically, on VSR and TextVQA, the 4-expert variant outperforms the 2-expert counterpart by 6.6% and 6.2%, respectively. Moreover, we also observe 1.2%, 1.1%, and 0.7% performance gains on HM, SciQA and IconQA. When the number of task experts becomes 8, it behaves similarly to the 4-expert case, which indicates that the benefit of increasing capacity converges as we use more task experts. Hence, we use 4 task experts as the default setting.
4.4.6 Temperature.

### v5

@@ 3.4 Universal Expert
As will be shown in Sec. 4.4.1, the formulation in Sec. 3.3 still hurts the generalization ability of the entire model, due to the fact that instruction-tuned models generalize to unseen tasks via training on extensive instructions Wei et al. (2021), while in our formulation, each expert sees fewer instructions than the original dense model.
To alleviate this problem, we propose an universal expert that learns from all training data. Specifically, we fix the number of activated experts to 1 (i.e., $k$ in Eq. 1 equals 1) and define $G_{\text{max}}$ as the maximum element in $\mathbf{G}$. Then the output for all the experts is expressed as:
@@
in which we additionally train an universal expert parameterized by $\mathbf{W}_{u}$. Different from the task experts that are activated only for specific model inputs, the universal expert is activated for all inputs. The final output is a weighted sum of outputs from one of the experts and the universal expert plus the original LVLM’s output. Consequently, the task expert learns distinct skills for certain tasks while the universal expert masters holistic understanding of the training corpus. The synergy between them offers both specialization and generalization for the LVLMs with MoCLE.
Models GQA VSR IQA Visdial MME POPE A-OKVQA OKVQA VQAv2 Direct MC (test) (test-dev) InstructBLIP (7B) 48.6 60.8 43.4 46.3 1202.9 77.6 58.8 73.8 57.0 77.4 + MoCLE 49.3 64.7 46.3 46.9 1222.6 82.1 61.5 78.2 59.8 78.9
Table 2: Zero-shot results (InstructBLIP) on the held-out datasets, i.e., GQA, VSR, IconQA (IQA), Visdial, MME, POPE and evaluation on held-in datasets, i.e., A-OKVQA, OKVQA, VQAv2. Here Direct and MC denote directly answering and multiple choices. Best results are marked in bold.
Models Flickr 30K VQAT HM SQA MSVD QA MSRVTT QA iVQA InstructBLIP (7B) 81.3 53.9 65.3 62.0 41.4 23.0 51.3 + MoCLE 81.9 57.1 65.6 63.9 42.6 24.4 53.2
Table 3: Zero-shot results (InstructBLIP) on the held-out datasets. Here, VQAT, HM and SQA denote TextVQA, HatefulMemes and ScienceQA, respectively.
Methods Train Data MME MMB SQA GeoQA VQA-RAD SLAKE PathVQA Open Closed Open Closed Open Closed Single LoRA LLaVA-665k 1804 65.89 67.67 - - - - - - - Geo170k - - - 57.82 - - - - - - Med. Mix - - - - 53.90 84.19 86.05 85.58 38.07 91.77 Single LoRA All 1794 64.69 66.78 57.56 46.89 77.94 84.61 82.45 35.56 90.71 MoCLE All 1838 66.07 67.38 60.21 53.59 81.98 83.29 85.10 35.21 91.65
Table 4: Evaluation results of LLaVA-1.5-7B, where MMB denotes MMBench.
@@ 4 Experiment
In this section, we conduct an assessment of MoCLE across multiple downstream tasks in a zero-shot setting. We first detail the experimental settings and implementation details, which are followed by a description of the datasets and instructions employed, along with the outcomes of our evaluations. Lastly, we present an ablation study and visualizations of clustering and routing results.
4.1 Implementation Details
We evaluate the effectiveness of MoCLE on two LVLMs: InstructBLIP Dai et al. (2023) and LLaVA-1.5 Liu et al. (2023a). Specifically, we compare the performance between the LVMs with and without MoCLE. The detailed configuration of MoCLE on these LVLMs are presented in Table 1. In addition, we encode all the instructions of different datasets using the all-MiniLM-L6-v2 variant of the Sentence Transformer model Reimers and Gurevych (2019) and cluster their embeddings via $k$-means clustering algorithm. More training details can be found in Appendix A.
4.2 Settings
For InstructBLIP, we follow Dai et al. (2023) for the choice of training datasets. However, these datasets only focus on a single domain: natural images. To validate the effectiveness of MoCLE on multiple domains, for LLaVA-1.5, in addition to its original training data LLaVA-665K Liu et al. (2023a) which focus on natural image domain, we include datasets from multiple domains, i.e., geometric tasks: Geo170k Gao et al. (2023a), medical tasks: VQA-RAD Lau et al. (2018), SLAKE Liu et al. (2021) and PathVQA He et al. (2020). More details on the training and evaluation datasets are provided in Appendix C. Note that during evaluation, we report the CIDEr score Vedantam et al. (2015) for Flickr30K, the iVQA accuracy for iVQA, AUC score for HatefulMemes, Mean Reciprocal Rank (MRR) for Visual Dialog, the perception/perception+cognition score for MME (InstructBLIP/LLaVA) and F1 score for the adversarial split of POPE. For all other datasets, we report the top-1 accuracy (%). Task templates for evaluation can be found in Appendix D.
LoRA(r=8) r=64 Cluster MoE Uni. Expert LoRA # Params. Flickr HM SQA IQA VSR VQAT Avg.
(a) 4.19M 81.3 65.1 57.4 44.2 62.8 49.4 60.0
[…]
Table 5: Comparison of individual components of the MoCLE framework in zero-shot vision-language tasks. Default settings are marked in gray.
@@ 4.3 Evaluation Results
InstructBLIP Tables 2 and 3 show the results on multiple held-out/in vision-language tasks. The proposed MoCLE shows considerable performance improvement over the original LVLM. Specifically, on held-out datasets such as IconQA, Visual Spatial Reasoning (VSR), TextVQA and ScienceQA datasets, we obtain an absolute performance gain of 2.9%, 3.9%, 3.2%, and 1.9%, respectively. On held-in datasets, an absolute improvement of 4.4%, 2.8% and 1.5% can be observed on A-OKVQA (MC), OKVQA and VQAv2, respectively. This indicates that the proposed MoCLE facilitates generalization to unseen tasks and can effectively alleviate task conflicts during multi-task learning
LLaVA-1.5 Similar to the preliminary results in Figure 1, we consider a Single-LoRA baseline where a single set of LoRAs are trained on natural images (LLaVA-665k), geometric (Geo170K), medical (Med. Mix) and a mixture of all tasks (All). As can be seen, due to task conflicts, the model trained on all tasks shows inferior results compared to those trained on only one task. However, MoCLE is able to reduce this gap on medical tasks and even offer better performance on natural image (MME, MMB) and geometric tasks (GeoQA) compared to the model trained on one task. This shows that MoCLE is effective with the presence of multiple-domain datasets.
@@ 4.4 Ablation Studies
In this section, we first ablate the effectiveness of the main components (i.e., Cluster MoE and universal expert) in the proposed MoCLE. Then we conduct a thorough analysis to study how the proposed MoCLE responds to changes in hyper-parameters (e.g., temperature and the number of clusters and task experts). Notice that we use InstructBLIP for all ablations and we report the evaluation results on Flickr30K (Flickr), Hateful Memes (HM), ScienceQA (SQA), IconQA (IQA), Visual Spatial Reasoning (VSR) and TextVQA (VQAT).
4.4.1 Effects of Different Components
We start from MoCLE and remove its key component one-by-one to analyze their effect.
Universal expert. Table 5 shows the ablation results when varying different components of MoCLE. By comparing rows (d) and (c), we notice that a sharp performance drop in VSR and TextVQA tasks when the universal expert is removed. This is due to that instruction-tuned model generalizes to unseen tasks by training on many instructions, while in our case, each expert sees fewer instructions than the dense model. For example, task TextVQA with instruction “OCR tokens: {}, Question: {}. Short answer:” needs not only VQA ability but also optical character recognition (OCR) skills, which are learned jointly from VQA data formatted as “Question: {}. Short answer:” and TextCaps data formatted as “OCR tokens: {}. Write a description for the photo”. Thus, universal expert is necessary to maintain generalization ability.
Cluster MoE. Comparing rows (c) and (a), we observe performance drop on SQA, IQA, and VQAT when cluster MoE is not used, which indicates it can alleviate task conflicts within a single set of LoRAs between different tasks.
LoRA rank. As can be seen from rows (c), (b) and (a), naïvely increasing the LoRA ranks from 8 to 64 only leads to a small average performance improvement of 0.7%, and thus cannot address task conflicts. Instead, promoting task specialization via clustering achieves notable improvement with fewer additional parameter (×$\times 4$ in Cluster MoE versus ×$\times 8$ when increasing the rank to 64).
4.4.2 Universal Expert vs. Top-2 Experts
To ablate the proposed universal expert, we remove it and activate one more existing expert. i.e., top-$2$ gating. We report their performance in Table 6. The top-2 MoE model yields inferior results compared to MoCLE and it even performs worse than the MoCLE variant without the universal expert reported in Table 5. This can be explained by the intensified conflicts when task experts are shared via top-$2$ gating because now each expert need to learn common feature with other experts. However, as the universal expert is shared all the time especially for this purpose, it frees task experts from this duty and thus alleviates the conflicts.
Flickr HM SQA IQA VSR VQAT Avg.
@@ 4.4.1 Effects of Different Components
Gating Flickr HM SQA IQA VSR VQAT Avg.
Token (LLaVA-MoLE) 81.7 65.4 61.9 44.0 49.0 46.6 58.1
Chen et al. (2024)
Sentence (Octavius) 82.0 65.1 62.3 45.3 56.6 47.0 59.7
Chen et al. (2023g)
Dataset 80.3 64.6 63.1 45.9 57.6 53.4 60.8
Jang et al. (2023)
Cluster 81.9 65.4 63.3 46.1 58.9 54.9 61.8
@@
Table 7: Ablation study on routing inputs based on different input conditions.
Figure 3: Ablation study on the number of clusters, experts and gate temperature. The x-axes of the first and last figures are log-scaled. The y-axes are the average performance of Flickr, HM, SQA, IQA, VSR and VQAT.
4.4.3 Gating Strategies
We compare the proposed cluster-conditioned gating strategy with existing MoLE methods in Table 7. Note that MoLE denotes the mixture of LoRA experts by applying MoE to LoRA. Some of these methods adopt different configurations, e.g.., # experts, expert params. and ranks. For fair comparison, we follow the first row of Table 1 except that universal expert is not enabled in this experiment.
@@ 4.4.3 Gating Strategies
Token/Sentence-MoLE. The former obtains the routing decision based on the hidden representations of each token (adopted by Chen et al. (2024)) and the later on the average representations of the instruction tokens while excluding the visual tokens. (adopted by Chen et al. (2023f)). Both of these methods give inferior results on the evaluation tasks. We speculate that this is because (1) a sparse expert learns less data than its dense counterpart, leading to lack of task generalization, (2) similar tasks are not grouped together by the same expert, resulting in task conflicts within that expert, which can be verified by the routing visualization in Figure 6, where samples in the same dataset are routed to multiple experts instead of a dedicated one.
Dataset-MoLE is a special case of MoCLE as it treats each dataset as a cluster while MoCLE leverages $k$-means to achieve this. It closely resembles to the dataset expert proposed in Jang et al. (2023) except that we assign clusters for a sentence by its distance to cluster centers while they count, to this sentence, the number of closest reference sentences belonging to each dataset. Further, for fair comparison, we only use 4 experts but they allocate an expert for each dataset. We observe inferior results compared to the proposed cluster routing. This results from the fact that Dataset-MoLE is less flexible as it can only assign a dataset to one cluster. However, in practice, we observe multiple tasks in a dataset which should be assigned to different clusters. (e.g., llava_150k contains reasoning, conversations and captioning, which are assigned to different clusters/experts as in Figure 5 and 6(a)).
4.4.4 Number of Clusters
@@
The number of clusters $K$ controls the granularity of task specialization. A very small $K$ would result in many different tasks to be processed by the same expert, and can increase the chance of task conflicts. As shown in Figure 3, when we cluster the inputs into 4 groups, the resulting model performs poorly on the evaluation tasks. However, as we increase the number of clusters to 16 and 64, we observe considerable performance gains. However, a $K$ too large (256) introduces unnecessary complexity to the routing process (e.g., a paraphrased instruction gets routed to different experts). So we use 64 clusters by default.
4.4.5 Temperature